### v2

@@ 4.2 Global iteration complexity
We now estimate the global iteration complexity of the RS-RNM under Assumption 4 and the following assumption.
@@
The following inequality can also be derived from $(iii)$:
$\left\|\nabla^{2}f(y)(y-x)-(\nabla f(x)-\nabla f(y))\right\|\leq\frac{L_{H}}{2}\left\|x-y\right\|^{2},\quad\forall x,y\in\Omega+B(0,r_{1}).$ (17)
From the definition of $r_{1}$ in $(iii)$, Lemma 4.1 and 9, we have
@@
Note that unlike (12), the bound has no dependency on $\varepsilon$. For this reason, we have
@@
$\left\|\nabla^{2}f(x)\right\|\leq U_{H},\ \forall x\in\Omega+B(0,r_{1}).$ (18)
@@
Suppose that Assumptions 4 and 4.2 hold. Then, with probability at least $1-2e^{-s}$, a step size $t_{k}^{\prime}>0$ such that
@@
Proof. As 14 is obtained in the proof of Lemma 4.5, there exists $\tau_{k}^{\prime}\in(0,1)$ such that
@@
Since we have $21-\alpha\geq 1/2\geq t^{\prime}_{k}/$ 2 from Assumption 4.2 $(ii)$, we obtain
@@
$-\frac{1}{2}{t_{k}^{\prime}}^{2}g_{k}^{\mathsf{T}}P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}(\nabla^{2}f(x_{k}+\tau_{k}^{\prime}t_{k}^{\prime}d_{k})-H_{k})P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}g_{k}.$ (19)
@@
Using Lemma 2.2 and Assumption 4.2 $(iii)$, we also obtain, with probability at least $1-2e^{-s}$, the bound of the second term:
@@
$=\frac{\bar{\mathcal{C}}n}{2s}L_{H}{t_{k}^{\prime}}^{2}\left\|d_{k}\right\|\left(\frac{c_{2}s\left\|g_{k}\right\|^{\gamma}}{\bar{\mathcal{C}}L_{H}n\left\|d_{k}\right\|}-t_{k}^{\prime}\right)\left\|M_{k}^{-1}P_{k}g_{k}\right\|^{2}.$
Moreover, from (9), (4.1) and Assumption 4.2 $(i)$, we have
@@
Suppose that Assumptions 4 and 4.2 hold. Then, with probability at least $1-2e^{-s}$, the step size $t_{k}$ chosen in Line 6 of RS-RNM satisfies
$t_{k}\geq t_{\min},$ (21)
@@
Since 21 is equivalent to $\beta^{l_{k}}\geq t_{\min}$, and moreover
@@
Suppose that Assumptions 4 and 4.2 hold. Let
[…]
Then, with probability at least
$1-2m\left(\exp(-\frac{{\mathcal{C}}_{0}}{4}s)-\exp(-s)\right),$ (22)
we have
$\sqrt{\frac{f(x_{0})-f^{*}}{mp}}\geq\min_{k=0,1,\dots,m-1}\left\|g_{k}\right\|.$
@@
Proof. Replacing $U_{H}(\varepsilon)$ and $t_{\min}(\varepsilon)$ with $U_{H}$, in (18), and $t_{\min}$ respectively in the argument in the proof of Theorem 4.9, we have
@@
with the given probability. Therefore, we obtain
@@
$\geq p\sum_{k=0}^{m-1}\left\|g_{k}\right\|^{2}$
$\geq mp\left(\min_{k=0,1,\dots,m-1}\left\|g_{k}\right\|\right)^{2},$
which is equivalent to the conclusion.
If we ignore the probability, Theorem 4.16 shows that we get $\left\|g_{k}\right\|\leq\varepsilon$ after at most $O(\varepsilon^{-2})$ iterations. This global complexity $O(\varepsilon^{-2})$ is the same as that obtained in [26] for the regularized Newton method.
4.3 Discussion on the probability bound
Let us now discuss the probability bound, (22), that appears in Theorem 4.16. In the following proposition, we prove that, under some basics assumptions, this probability is in the order of $1-o(1)$ by taking $s=O(\log^{2}(n))$.
Proposition 4.18.
Assuming that $f$, $\nabla f$ and $\nabla^{2}f$ are respectively $L_{f}$, $L_{g}$ and $L_{H}$-Lipschitz on $\Omega$ and that the value $\sup_{k}\|x_{k}-{x}^{*}\|_{\infty}$, where ${x}^{*}$ denotes an optimal solution, is independent of the dimension $n$. Then by choosing $s=O(\log^{2}(n))$, we have that the probability of success of Theorem 4.16 tends to $1$ as $n$ tends to infinity.
Proof 4.19.
Proof. Notice that since $f$ is Lipschitz, and that $\|x_{0}-x^{*}\|\leq\|x_{0}-x^{*}\|_{\infty}\sqrt{n}$, we have that
$f(x_{0})-f^{*}\leq L_{f}\|x_{0}-x^{*}\|_{\infty}\sqrt{n}$
for some constant. Hence assuming that $\|x_{0}-x^{*}\|_{\infty}$ is independent of the ambient dimension $n$, we deduce that
$f(x_{0})-f^{*}\leq\hat{L}\sqrt{n},$
for some constant $\hat{L}$. Similarly since
$p=\frac{\alpha t_{\min}}{2\bar{\mathcal{C}}(1+c_{1})\frac{n}{s}U_{H}+2c_{2}U_{g}^{\gamma}}\quad\&\quad t_{\min}=\min\left(1,\frac{\beta c_{2}^{2}s^{2}}{\bar{\mathcal{C}}^{2}L_{H}U_{g}^{1-2\gamma}n^{2}}\right),$
we have that
$\frac{1}{p}\leq\max\left(1,\frac{\bar{\mathcal{C}}^{2}L_{H}U_{g}^{1-2\gamma}n^{2}}{\beta c_{2}^{2}s^{2}}\right)\frac{2\bar{\mathcal{C}}(1+c_{1})\frac{n}{s}U_{H}+2c_{2}U_{g}^{\gamma}}{\alpha}.$
Since $\nabla f(x)$ is $L_{g}$-Lipschitz and $\bar{x}$ is a stationary point, we have that $\|\nabla f(x_{k})\|\leq L_{g}\|x_{k}-\bar{x}\|_{2}\leq L_{g}\sqrt{n}\|x_{k}-\bar{x}\|_{\infty}$.
Furthermore we also have that $\|\nabla^{2}f(x_{k})\|\leq\|\nabla^{2}f(\bar{x})\|+L_{H}\sqrt{n}\|x_{k}-\bar{x}\|_{\infty}$. Hence $U_{g}$ in (9) and $U_{H}$ in (18) can be bounded by $O(\sqrt{n})$. We can therefore bound $m$ by $\hat{C}\frac{n^{4}}{\varepsilon^{2}}$ for some constant $\hat{C}$ large enough. Hence, by taking $s=\tilde{C}\log^{2}(n)$ for any constant $\tilde{C}$, we see that for such $s$, the term $2m\left(\exp(-\frac{{\mathcal{C}}_{0}}{4}s)-\exp(-s)\right)$ in (22) tends to 00 as $n$ tends to infinity.
Notice that, by a cubic regularization, the R-ARC algorithm in [24] achieves $O(\varepsilon^{-3/2})$ to obtain a first order stationary point.
@@ 5 Local convergence
In this section, we investigate local convergence properties of the sequence $\{x_{k}\}$ assuming that it converges to a strict local minimal $\bar{x}$. First we will show that the sequence converges locally linearly to the strict local minimizer; then we will prove that, when $f$ is strongly convex, we cannot aim at local super-linear convergence using random subspace. Finally, we will prove that when the Hessian at $\bar{x}$ is rank deficient then we can attain super-linear convergence for $s<n$ large enough.
@@ 5.1 Local linear convergence
In this subsection we will show the sequence $\{f(x_{k})-f(\bar{x})\}$ converges locally linearly, i.e. there exists $\kappa\in(0,1)$ such that for $k$ large enough,
@@
We will further prove that $\kappa$ can be expressed as $\kappa=O(\frac{s}{n\tilde{\kappa}(\nabla^{2}f(\bar{x}))})$, where $\tilde{\kappa}(\nabla^{2}f(\bar{x}))$ is the ratio of the largest eigenvalue value over the smallest non-zero eigenvalue of $\nabla^{2}f(\bar{x})$. Notice that, to the best of our knowledge, until now, local linear convergence is always proved for subspace algorithms assuming that the function is locally strongly convex or satisfies some PL-inequality (3). In this section we prove that under a Hölderian error bound condition, and an additional mild assumptions on the rank of the Hessian at the local minimizer, we can prove local linear convergence. More precisely let us denote by $r=\mathrm{rank}(\nabla^{2}f(\bar{x}))$, which measures the number of positive eigenvalues of $\nabla^{2}f(\bar{x})$. We will first prove that, for any $\sigma\in(0,1)$, there exists $n_{0}$ such that if $n\geq n_{0}$ and $r\geq\sigma n$, then for any $x$ in the a neighborhood of $\bar{x}$, the function
$\tilde{f}_{x}:u\mapsto f(x+P^{\top}u)$
is strongly convex with high probability in a neighborhood of 00. Let us fix $\sigma\in(0,1)$. We recall here that $P\in\mathbb{R}^{s\times n}$ is equal to $\frac{1}{\sqrt{s}}$ times a random Gaussian matrix. In this subsection, we make the following additional assumptions:
{assumption}
We have that $s=o(n)$, that is, $\lim\limits_{n\rightarrow+\infty}\frac{s}{n}=0$.
@@
From Assumption 5.1 $(i)$, $\nabla^{2}f(\bar{x})$ has $r$ positive eigenvalues, i.e, $\lambda_{1}(\bar{x})\geq\cdots\lambda_{r}(\bar{x})>0$. By continuity of the eigenvalues, there exists a neighborhood $\bar{B}$ of $\bar{x}$ such that for any $x\in\bar{B}$, $\lambda_{r}(x)\geq\frac{\lambda_{r}(\bar{x})}{2}$. Here, we assume, w.l.o.g. that $\bar{B}\subseteq B_{H}$, where $B_{H}$ is defined in Assumption 5. Let us denote
$\bar{\lambda}:=\frac{\lambda_{r}(\bar{x})}{2}.$ (23)
Assumption 5.1 $(ii)$ is called a Hölderian growth condition or a Hölderian error bound condition [14]. The condition is weaker than local strong convexity in the sense that it holds with $\rho=2$ if $f$ is locally strongly convex.
@@
Let $10<\varepsilon_{0}<$ 1. Then under Assumptions 5.1 and 5.1 $(i)$, there exist $n_{0}\in\mathbb{N}$ (which depends only on $\varepsilon_{0}$ and $\sigma$) and a neighborhood $B^{*}\subseteq\bar{B}$ such that if $n\geq n_{0}$, for any $x\in B^{*}$,
$\nabla^{2}\tilde{f}_{x}(0)\succeq\frac{(1-\varepsilon_{0})^{2}n}{2s}\sigma\bar{\lambda}I_{s}$
@@
Proof. Let $x\in\bar{B}$ be fixed. Because of $\nabla^{2}\tilde{f}_{x}(0)=P\nabla^{2}f(x)P^{\top}$, we have $u^{\top}\nabla^{2}\tilde{f}_{x}(0)u=(P^{\top}u)^{\top}\nabla^{2}f(x)(P^{\top}u)$ for any $u\in\mathbb{R}^{s}$. Let $\nabla^{2}f(x)=U(x)D(x)U(x)^{\top}$ be the eigenvalue decomposition of $\nabla^{2}f(x)$. Since $\nabla^{2}\tilde{f}_{x}(0)=(PU(x))D(x)(PU(x))^{\top}$ and $PU(x)$ has the same distribution as $P$, we can assume here w.l.o.g. that $PU(x)=P$. Here
@@
$(P^{\top}u)^{\top}D(x)(P^{\top}u)=(P^{1}u)^{\top}D_{1}(x)(P^{1}u)+(P^{2}u)^{\top}D_{2}(x)(P^{2}u).$ (24)
By Assumption 5.1 $(i)$, and by definition of $\bar{B}$, we have that $D_{1}(x)\succeq\lambda_{r}(x)I_{n_{1}}\succeq\bar{\lambda}I_{n_{1}}\succ 0$, and $D_{2}(x)\succeq\lambda_{n}(x)I_{n_{2}}$. Hence from (24), we have
$(P^{\top}u)^{\top}D(x)(P^{\top}u)\geq\bar{\lambda}\|P^{1}u\|^{2}+\lambda_{n}(x)\|P^{2}u\|^{2}.$ (25)
Let $\sigma_{\max}(\cdot)$ and $\sigma_{\min}(\cdot)$ denote respectively the largest and the smallest singular value of a matrix. Using [28, Theorem 4.6.1], there exists a constant $\mathcal{C}$ such that with probability at least $1-6\exp(-s)$:
$\sqrt{\frac{n}{s}}-\mathcal{C}\leq\sigma_{\min}(P^{\top})\leq\sigma_{\max}(P^{\top})\leq\sqrt{\frac{n}{s}}+\mathcal{C},$ (26)
@@
More precisely, since all the three matrices $P^{\top},P^{1}$ and $P^{2}$ are Gaussian random matrices, we can apply [28, Theorem 4.6.1] and deduce that each of the three inequalities above holds with probability $1-2\exp(-s)$. The probability that all the three equations hold is derived using (4). Hence, with probability at least $1-6e^{-s}$, for any $u\in\mathbb{R}^{s}$,
@@
We have that $n_{1}\geq\sigma n$ and $n_{2}\leq(1-\sigma)n$. Since $s=o(n)$ by Assumption 5.1, we also have $s=o(n_{1})$. Hence
$\lim_{n\to\infty}\frac{\sqrt{\frac{n_{1}}{s}}-\mathcal{C}}{\sqrt{n/s}}\geq\sqrt{\sigma}\quad\&\quad\lim_{n\to\infty}\frac{\sqrt{\frac{n_{2}}{s}}+\mathcal{C}}{\sqrt{n/s}}\leq\sqrt{(1-\sigma)}.$
Let $10<\varepsilon_{0}<$ 1, there exists $n_{0}$ such that if $n\geq n_{0}$,
$\|P^{1}u\|\geq$ $(1-\varepsilon_{0})\sqrt{\sigma(n/s)}\|u\|,$
$\|P^{2}u\|\leq$ $\varepsilon_{0}\sqrt{(1-\sigma)(n/s)}\|u\|.$
Hence, from (25), we have that
$(P^{\top}u)^{\top}D(x)(P^{\top}u)\geq{n/s}\left((1-\varepsilon_{0})^{2}\sigma\bar{\lambda}+\varepsilon_{0}^{2}(1-\sigma)\min(\lambda_{n}(x),0)\right)\|u\|^{2}.$
We conclude the proposition by noticing that $\min(\lambda_{n}(x),0)$ tends to 00, hence the claim holds by considering a neighborhood $B^{*}\subseteq\bar{B}$ of $\bar{x}$ small enough.
@@
Assume that Assumptions 5.1 and 5.1 $(i)$ hold, there exist $n_{0}\in\mathbb{N}$ (which depends only on $\varepsilon_{0}$ and $\sigma$) and neighborhoods $\hat{B}\subset B^{*}$ and $B_{0}$ (a neighborhood of $0\in\mathbb{R}^{s}$) such that if $n\geq n_{0}$, for any $x\in\hat{B}$,
${\nabla\tilde{f}_{x}(0)^{\top}(P_{k}\nabla^{2}f({x})P_{k}^{\top})^{-1}\nabla\tilde{f}_{x}(0)\geq f(x)-\min\limits_{u\in B_{0}}f(x+P^{\top}u)}$
@@
Proof. Let $\hat{B}\subset B^{*}$, and let $x\in\hat{B}$. By the Taylor expansion of $\tilde{f}_{x}$ at 00 we have by Proposition 5.1 that for any $u\in\mathbb{R}^{s}$ such that $x+P^{\top}u\in B^{*}$,
$f(x+P^{\top}u)\geq f(x)+(P\nabla f(x))^{\top}u+\frac{1}{4}u^{\top}P_{k}\nabla^{2}f({x})P_{k}^{\top}u.$
Let $B_{0}$ be a neighborhood of $0\in\mathbb{R}^{s}$ such that $x+P^{\top}u\in B^{*}$ for any $x\in\hat{B}$. Let $g(u)=(P\nabla f(x))^{\top}u+\frac{1}{4}u^{\top}P_{k}\nabla^{2}f({x})P_{k}^{\top}u$. By the above inequality we have that
$\min\limits_{u\in B_{0}}f(x+P^{\top}u)\geq f(x)+\min\limits_{u\in B_{0}}g(u).$ (27)
By Proposition 5.1 we know that for any $u\in\mathbb{R}^{s}$ such that $x+P^{\top}u\in B^{*}$, $g$ is convex, and the minimum is attained at the point $u^{*}$ satisfying
$\nabla g(u^{*})=P\nabla f(x)+\frac{1}{2}P_{k}\nabla^{2}f({x})P_{k}^{\top}u^{*}=0.$
Hence, since $\|\nabla f(x)\|$ tends to 00 as $x$ tends to $\bar{x}$, we can ensure, by taking $\hat{B}$ small enough, that $u^{*}\in B_{0}$. Hence
$\min\limits_{u\in B_{0}}g(u)$ $=-{2}(P_{k}\nabla f(x))^{\top}(P_{k}\nabla^{2}f({x})P_{k}^{\top})^{-1}P_{k}\nabla f(x)+\frac{1}{4}{4}(P_{k}\nabla f(x))^{\top}(P_{k}\nabla^{2}f({x})P_{k}^{\top})^{-1}P_{k}\nabla f(x)$
$=-(P_{k}\nabla f(x))^{\top}(P_{k}\nabla^{2}f({x})P_{k}^{\top})^{-1}P_{k}\nabla f(x)$
holds and (27) yields the desired inequality.
@@
Assume that Assumptions 5, 5.1, and 5.1 hold. There exists $k_{0}\in\mathbb{N}$ and $n_{0}\in\mathbb{N}$ such that if $n\geq n_{0}$ and $k\geq k_{0}$, we have with probability $1-6(\exp(-s)+\exp(-\frac{\mathcal{C}_{0}}{4}s))$:

### v3

@@ 4.2 Global iteration complexity
We now estimate the global iteration complexity of the RS-RNM under Section 4 and the following assumption.
@@
From the definition of $r_{1}$ in $(iii)$, Lemma 4.1 and Eq. 9, we have
@@
Note that unlike Eq. 12, the bound has no dependency on $\varepsilon$. For this reason, we have
@@
(20) $\left\|\nabla^{2}f(x)\right\|\leq U_{H},\ \forall x\in\Omega+B(0,r_{1}).$
@@
Suppose that Section 4 and Section 4.2 hold. Then, with probability at least $1-2e^{-s}$, a step size $t_{k}^{\prime}>0$ such that
@@
Proof. As Eq. 14 is obtained in the proof of Lemma 4.5, there exists $\tau_{k}^{\prime}\in(0,1)$ such that
@@
Since we have $1-\alpha\geq 1/2\geq t^{\prime}_{k}/2$ from Section 4.2 $(ii)$, we obtain
@@
[…]
@@
Using Lemma 2.2 and Section 4.2 $(iii)$, we also obtain, with probability at least $1-2e^{-s}$, the bound of the second term:
@@
(23) $=\frac{\bar{\mathcal{C}}n}{2s}L_{H}{t_{k}^{\prime}}^{2}\left\|d_{k}\right\|\left(\frac{c_{2}s\left\|g_{k}\right\|^{\gamma}}{\bar{\mathcal{C}}L_{H}n\left\|d_{k}\right\|}-t_{k}^{\prime}\right)\left\|M_{k}^{-1}P_{k}g_{k}\right\|^{2}.$
Moreover, from Eq. 9, Lemma 4.1 and Section 4.2 $(i)$, we have
@@
Suppose that Section 4 and Section 4.2 hold. Then, with probability at least $1-2e^{-s}$, the step size $t_{k}$ chosen in Line 6 of RS-RNM satisfies
(24) $t_{k}\geq t_{\min},$
@@
Since Eq. 24 is equivalent to $\beta^{l_{k}}\geq t_{\min}$, and moreover
@@
Suppose that Section 4 and Section 4.2 hold. Consider any $\delta\in(0,1)$. Let
$m=\left\lfloor\frac{f(x_{0})-f^{*}}{(1-\delta)(1-\delta_{s})p\varepsilon^{2}}\right\rfloor+1,\quad\text{where}\quad p=\frac{\alpha t_{\min}}{2\bar{\mathcal{C}}(1+c_{1})\frac{n}{s}U_{H}+2c_{2}U_{g}^{\gamma}},$
and where $\delta_{s}=2\left(\exp(-\frac{{\mathcal{C}}_{0}}{4}s)-\exp(-s)\right)$. Then, we have that
$\sqrt{\frac{f(x_{0})-f^{*}}{mp}}\geq\min_{k=0,1,\dots,m-1}\left\|g_{k}\right\|$
holds with probability at least $1-\exp\left(-\frac{\delta^{2}}{2}(1-\delta_{s})m\right)$.
@@
Proof. Replacing $U_{H}(\varepsilon)$ and $t_{\min(\varepsilon)}$ with $U_{H}$, in Eq. 20, and $t_{\min}$ respectively in the argument in the proof of Theorem 4.9, we have
@@
with the given probability. Therefore, by using the same notation as in the proof of Theorem 4.9, we obtain:
@@
$\geq p\sum_{k=0}^{m-1}\left\|g_{k}\right\|^{2}T_{k}$
$\geq p\left(\min_{k=0,1,\dots,m-1}\left\|g_{k}\right\|^{2}\right)\sum_{k=0}^{m-1}T_{k}$
$\geq(1-\delta)(1-\delta_{s})mp\left(\min_{k=0,1,\dots,m-1}\left\|g_{k}\right\|^{2}\right),$
where the last inequality holds with probability $1-\exp\left(-\frac{\delta^{2}}{2}(1-\delta_{s})m\right)$ as shown in (18). This prove the theorem.
If we ignore the probability, Theorem 4.16 shows that we get $\left\|g_{k}\right\|\leq\varepsilon$ after at most $O(\varepsilon^{-2})$ iterations. This global complexity $O(\varepsilon^{-2})$ is the same as that obtained in [39] for the regularized Newton method. Notice that, by a cubic regularization, the R-ARC algorithm in [37] achieves $O(\varepsilon^{-3/2})$ to obtain a first order stationary point.
@@ 5 Local convergence
In this section, we investigate local convergence properties of the sequence $\{x_{k}\}$ assuming that it converges to a strict local minimizer $\bar{x}$. First we will show that the sequence converges locally linearly to the strict local minimizer; then we will prove that, when $f$ is strongly convex, we cannot aim at local super-linear convergence using random subspace. Finally, we will prove that when the Hessian at $\bar{x}$ is rank deficient then we can attain super-linear convergence for $s<n$ large enough.
@@ 5.1 Local linear convergence
In this subsection we will show that the sequence $\{f(x_{k})-f(\bar{x})\}$ converges locally linearly, i.e. there exists $\kappa\in(0,1)$ such that for $k$ large enough,
@@
We will further prove that $\kappa$ can be expressed as $\kappa=O(\frac{s}{n\tilde{\kappa}(\nabla^{2}f(\bar{x}))})$, where $\tilde{\kappa}(\nabla^{2}f(\bar{x}))$ is the ratio of the largest eigenvalue value over the smallest non-zero eigenvalue of $\nabla^{2}f(\bar{x})$. Notice that, to the best of our knowledge, until now, local linear convergence is always proved for subspace algorithms assuming that the function is locally strongly convex or satisfies some PL-inequality (3). In this section we prove that under a Hölderian error bound condition, and an additional mild assumptions on the rank of the Hessian at the local minimizer, we can prove local linear convergence. More precisely let us denote by $r=\mathrm{rank}(\nabla^{2}f(\bar{x}))$, which measures the number of positive eigenvalues of $\nabla^{2}f(\bar{x})$. We will first prove, under some assumption on the rank of the Hessian at $\bar{x}$ and on $s$, that for any $x$ in the a neighborhood of $\bar{x}$, the function
[…]
is strongly convex with high probability in a neighborhood of 0. Let us fix $\sigma\in(0,1)$. We recall here that $P\in\mathbb{R}^{s\times n}$ is equal to $\frac{1}{\sqrt{s}}$ times a random Gaussian matrix. In this subsection, we make the following additional assumptions:
@@
{assumption}
We have that $s\leq\min\left(\frac{\sigma}{4\mathcal{C}^{2}},\frac{4(1-\sigma)}{\mathcal{C}^{2}}\right)n$.
From Section 5.1 $(i)$, $\nabla^{2}f(\bar{x})$ has $r$ positive eigenvalues, i.e, $\lambda_{1}(\bar{x})\geq\cdots\lambda_{r}(\bar{x})>0$. By continuity of the eigenvalues, there exists a neighborhood $\bar{B}$ of $\bar{x}$ such that for any $x\in\bar{B}$, $\lambda_{r}(x)\geq\frac{\lambda_{r}(\bar{x})}{2}$. Here, we assume, w.l.o.g. that $\bar{B}\subseteq B_{H}$, where $B_{H}$ is defined in Section 5. Let us denote
(26) $\bar{\lambda}:=\frac{\lambda_{r}(\bar{x})}{2}.$
Section 5.1 $(ii)$ is called a Hölderian growth condition or a Hölderian error bound condition [24]. The condition is weaker than local strong convexity in the sense that it holds with $\rho=2$ if $f$ is locally strongly convex.
@@
Assume that Section 5.1 $(i)$ and Section 5.1 hold. Let us consider $\tilde{f}_{x}$ defined by (25). There exists a neighborhood $B^{*}\subseteq\bar{B}$ such that for any $x\in B^{*}$,
$\nabla^{2}\tilde{f}_{x}(0)\succeq\frac{n}{8s}\sigma\bar{\lambda}I_{s}$
@@
Proof. Let $x\in\bar{B}$ be fixed and let $P\in\mathbb{R}^{s\times n}$ be a Gaussian matrix. Because of $\nabla^{2}\tilde{f}_{x}(0)=P\nabla^{2}f(x)P^{\top}$, we have $u^{\top}\nabla^{2}\tilde{f}_{x}(0)u=(P^{\top}u)^{\top}\nabla^{2}f(x)(P^{\top}u)$ for any $u\in\mathbb{R}^{s}$. Let $\nabla^{2}f(x)=U(x)D(x)U(x)^{\top}$ be the eigenvalue decomposition of $\nabla^{2}f(x)$. Since $\nabla^{2}\tilde{f}_{x}(0)=(PU(x))D(x)(PU(x))^{\top}$ and $PU(x)$ has the same distribution as $P$, we can assume here w.l.o.g. that $PU(x)=P$. Here
@@
(27) $(P^{\top}u)^{\top}D(x)(P^{\top}u)=(P^{1}u)^{\top}D_{1}(x)(P^{1}u)+(P^{2}u)^{\top}D_{2}(x)(P^{2}u).$
By Section 5.1 $(i)$, and by definition of $\bar{B}$, we have that $D_{1}(x)\succeq\lambda_{r}(x)I_{n_{1}}\succeq\bar{\lambda}I_{n_{1}}\succ 0$, and $D_{2}(x)\succeq\lambda_{n}(x)I_{n_{2}}$. Hence from Eq. 27, we have
(28) $(P^{\top}u)^{\top}D(x)(P^{\top}u)\geq\bar{\lambda}\|P^{1}u\|^{2}+\lambda_{n}(x)\|P^{2}u\|^{2}.$
Let $\sigma_{\max}(\cdot)$ and $\sigma_{\min}(\cdot)$ denote respectively the largest and the smallest singular value of a matrix. Using [41, Theorem 4.6.1], there exists a constant $\mathcal{C}$ such that with probability at least $1-6\exp(-s)$:
(29) $\sqrt{\frac{n}{s}}-\mathcal{C}\leq\sigma_{\min}(P^{\top})\leq\sigma_{\max}(P^{\top})\leq\sqrt{\frac{n}{s}}+\mathcal{C},$
@@
More precisely, since all the three matrices $P^{\top},P^{1}$ and $P^{2}$ are Gaussian random matrices, we can apply [41, Theorem 4.6.1] and deduce that each of the three inequalities above holds with probability $1-2\exp(-s)$. The probability that all the three equations hold is derived using Eq. 4. Hence, with probability at least $1-6e^{-s}$, for any $u\in\mathbb{R}^{s}$,
@@
We have that $n_{1}\geq\sigma n$ and $n_{2}\leq(1-\sigma)n$. Furthermore, we have by Section 5.1 that $s\leq\frac{\sigma}{4\mathcal{C}^{2}}n$ implies that $\sqrt{\frac{\sigma n}{s}}-\mathcal{C}\geq\frac{1}{2}\sqrt{\frac{\sigma n}{s}}$ and $s\leq\frac{4(1-\sigma)}{4\mathcal{C}^{2}}n$ implies that $\sqrt{\frac{(1-\sigma)n}{s}}+\mathcal{C}\leq 2\sqrt{\frac{(1-\sigma)n}{s}}$. Hence
$\frac{\sqrt{\frac{n_{1}}{s}}-\mathcal{C}}{\sqrt{n/s}}\geq\frac{1}{2}\sqrt{\sigma}\quad\&\quad\frac{\sqrt{\frac{n_{2}}{s}}+\mathcal{C}}{\sqrt{n/s}}\leq 2\sqrt{(1-\sigma)}.$
Therefore,
$\|P^{1}u\|\geq$ ${\frac{1}{2}}\sqrt{\sigma(n/s)}\|u\|,$
$\|P^{2}u\|\leq$ ${2}\sqrt{(1-\sigma)(n/s)}\|u\|.$
Hence, from Eq. 28, we have that
$(P^{\top}u)^{\top}D(x)(P^{\top}u)\geq{n/s}\left(\frac{1}{{4}}\sigma\bar{\lambda}+{4}(1-\sigma)\min(\lambda_{n}(x),0)\right)\|u\|^{2}.$
We conclude the proposition by noticing that $\min(\lambda_{n}(x),0)$ tends to 0, hence the claim holds by considering a neighborhood $B^{*}\subseteq\bar{B}$ of $\bar{x}$ small enough.
@@
Assume that Section 5, Section 5.1 $(i)$ and Section 5.1 hold, and let $P\in\mathbb{R}^{s\times n}$ be a Gaussian matrix. There exist neighborhoods $\hat{B}\subset B^{*}$ and $B_{0}$ (a neighborhood of $0\in\mathbb{R}^{s}$) such that for any $x\in\hat{B}$,
$\nabla\tilde{f}_{x}(0)^{\top}(P\nabla^{2}f({x})P^{\top})^{-1}\nabla\tilde{f}_{x}(0)\geq f(x)-\min\limits_{u\in B_{0}}f(x+P^{\top}u)$}
@@
Proof. Let $\hat{B}\subset B^{*}$, and let $x\in\hat{B}$. By the Taylor expansion of $\tilde{f}_{x}$ at 0, there exists $\tilde{x}\in[x,x+P^{\top}u]$ such that
$f(x+P^{\top}u)=f(x)+(P\nabla f(x))^{\top}u+\frac{1}{2}u^{\top}P\nabla^{2}f({\tilde{x}})P^{\top}u.$
Since, by Proposition 5.1, we have that $P\nabla^{2}f(\tilde{x})P^{\top}\succ 0$ for any $x+P^{\top}u\in B^{*}$, we deduce by Section 5 that for $u$ small enough:
(30) $f(x+P^{\top}u)\geq f(x)+(P\nabla f(x))^{\top}u+\frac{1}{4}u^{\top}P\nabla^{2}f({x})P^{\top}u.$
Let $B_{0}$ be a neighborhood of $0\in\mathbb{R}^{s}$ such that, (30) holds, and $x+P^{\top}u\in B^{*}$ for any $x\in\hat{B}$. Let $g(u)=(P\nabla f(x))^{\top}u+\frac{1}{4}u^{\top}P\nabla^{2}f({x})P^{\top}u$. By the above inequality we have that
(31) $\min\limits_{u\in B_{0}}f(x+P^{\top}u)\geq f(x)+\min\limits_{u\in B_{0}}g(u).$
By Proposition 5.1 we know that for any $u\in\mathbb{R}^{s}$ such that $x+P^{\top}u\in B^{*}$, $g$ is convex. Thus, the minimum is attained at the point $u^{*}$ satisfying
$\nabla g(u^{*})=P\nabla f(x)+\frac{1}{2}P\nabla^{2}f({x})P^{\top}u^{*}=0.$
Hence, since $\|\nabla f(x)\|$ tends to 0 as $x$ tends to $\bar{x}$, we can ensure, by taking $\hat{B}$ small enough, that $u^{*}\in B_{0}$. Hence
$\min\limits_{u\in B_{0}}g(u)$ $=-{2}(P\nabla f(x))^{\top}(P\nabla^{2}f({x})P^{\top})^{-1}P\nabla f(x)+\frac{1}{4}{4}(P\nabla f(x))^{\top}(P\nabla^{2}f({x})P^{\top})^{-1}P\nabla f(x)$
$=-(P\nabla f(x))^{\top}(P\nabla^{2}f({x})P^{\top})^{-1}P\nabla f(x)$
holds and Eq. 31 yields the desired inequality.
@@
Assume that Section 5, Section 5.1, and Section 5.1 hold. There exists $k_{0}\in\mathbb{N}$ such that if $k\geq k_{0}$, we have with probability $1-6(\exp(-s)+\exp(-\frac{\mathcal{C}_{0}}{4}s))$: